solve 8/x - 5/y = 34, 3/x - 2/y = 13, where x not = 0, and y not = 0
Answer by Alan3354(69443)

solve 8/x - 5/y = 34, 3/x - 2/y = 13, where x not = 0, and y not = 0
----------
Sub r = 1/x, s = 1/y
----
8r - 5s = 34
3r - 2s = 13
------
16r - 10s = 68
15r - 10s = 65
------------------ Subtract
r = 3
-----
Find s
Sub to get x and y